# QCD phase transition at finite temperature and chemical potential with the non-extensive statistics

### v1

@@ Abstract
With the application of the non-extensive correction to the equation of state (EoS) in the parton (hadron resonance) gas at high (low) temperature and a smooth interpolation of these two equation of states, a non-extensive version of EoS is constructed to investigate the properties of quantum chromodynamics (QCD) phase transition at finite temperature ($T$) and chemical potential ($\mu$). It is based on non-extensive statistics and characterized by a non-extensivity parameter $q$, which measures the degrees of deviation from the thermal equilibrium. We explore the QCD phase transition with $\mu=0$ MeV in the chiral and non-chiral limit cases, as well as the transition with finite $\mu$ in the non-chiral limit case. It is found that the dimensionless thermodynamic quantities such as the entropy density, the pressure, the energy density, the specific heat at constant volume and the trace anomaly are sensitive to the deviation of $q$ from unity and they become large both in the hadronic and quark-gluon plasma (QGP) phases with the increase of $q$. In the chiral limit case, the sound velocity squared ($(c_{s}^{2})_{q}$) is insensitive to the deviation. However, in the non-chiral limit case, this deviation leads to nontrivial corrections of $(c_{s}^{2})_{q}$ in the vicinity of the critical point ($T_{c}$) and at lower temperatures. Moreover, these dimensionless thermodynamic quantities are sensitive to the deviation of $\mu$ from zero and they become large both in the hadronic and QGP phases with the increase of $\mu$. However, for $(c_{s}^{2})_{q}$, it increases in the vicinity of $T_{c}$ and decreases at lower temperature. Finally, we observe that our results in the non-chiral limit case with $q=1$ agree well with those from the Lattice QCD, the hadron resonance gas model, and the Thermal-Fist fit to the hadron yields in high energy nuclear collisions in the low temperature region up to $T\sim 150$ MeV.
@@ I Introduction
In recent years, several effective models, such as the bag model [1], the Nambu-Jona-Lasinio (NJL) model [2, 3], the linear sigma model [4], the quark-meson model [5, 6], the Dyson-Schwinger equations approach [7, 8], the functional renormalization group approach [9, 10, 11], the hadron resonance gas (HRG) and statistical models [12, 13, 14, 15], and the Lattice QCD (LQCD) [14, 15, 16, 17], have been utilized to investigate the properties of the QCD phase transition. In these models, a thermal equilibrium is assumed and the extensive statistical approach, such as the Boltzmann-Gibbs (BG) statistics, the Bose-Einstein statistics or the Fermi-Dirac statistics is applied.
@@
However, such an approach is only valid when the heat bath of the system is homogeneous. In reality, this condition cannot always be satisfied. For example, in high-energy nucleus-nucleus (AA) collisions, extreme conditions of density and temperature give rise to strong intrinsic fluctuations and long-range correlations [18, 19, 20, 21]. Thus, instead of an equilibrium state, some kind of stationary state near the equilibrium ($q$-equilibrium) is expected to be formed [22, 23]. This phenomenon can be dealt with non-extensive statistics (Tsallis statistics) [24] where the degree of deviation from the equilibrium is described by a single parameter $q$.
@@
In recent years, Tsallis statistics has been widely employed to describe the multi-particle production processes in high-energy proton-proton (pp) [25, 26, 27, 28, 29, 30, 31, 32, 33] and AA [34, 35, 36, 37, 38, 39, 40, 41, 42] collisions. In most of these applications, the Tsallis distribution or the Tsallis-extended blast-wave model is simply fitted to the particle transverse momentum ($p_{\rm T}$) spectra. The extracted values of $q$ from the fits are usually larger than 1. The deviation of $q$ from unity represents the intrinsic fluctuations of the temperature [18] or of the mean value of the charged-particle multiplicity [43] in the hadronizing system. In case of high-energy AA collisions, the fluid dynamics is not considered in these applications, which could affect the accuracy of the results. In light of such a situation, (0+1)-, (1+1)-, (2+1)- and (3+1)-dimensional hydrodynamic models based on non-extensive statistics were developed as a description of relativistic heavy-ion collisions [44, 45, 46, 47]. This is a reasonable extension of the hydrodynamic model, as far-from-equilibrium evolution can still obey the hydrodynamic equations of motion, provided that the system is locally isotropized in the co-moving frame of the fluid [48, 49]. It was shown that the non-extensive hydrodynamic model can describe the charged hadron $p_{\rm T}$ spectra up to 6-8 GeV/c, which is larger than the typical validity range of the extensive hydrodynamic description, 2-3 GeV/c. Moreover, the elliptic flow of charged particles was reduced with the introduction of Tsallis statistics to the hydrodynamic model. The Tsallis statistics has also been applied on the nuclear equation of state (EoS) and the QGP EoS in the framework of a relativistic mean-field approach and the bag model, respectively. In the former model, the interactions of nucleons are mediated by the exchange of isoscalar and isovector mesons, $\sigma$, $\omega$ and $\rho$ [50, 51, 52, 53]. In the latter model, the quark matter is described as a gas of free quarks and the non-perturbative effects are simulated by the bag constant $B$ [50, 51, 53]. It was observed that both the nuclear EoS and the QGP EoS became stiffer when the non-extensive statistic effect was considered. Furthermore, the Tsallis statistics has been taken into account in the NJL model to describe the QCD phase transition [54, 55, 56, 57]. It was found that with the increase of $q$ the critical end point moved toward the direction of larger chemical potential and lower temperature.
In this study, we try to construct a non-extensive version of EoS to investigate the generic feature of the QCD phase transition at finite temperature and chemical potential. We first explore the QCD phase transition with zero chemical potential in the chiral and non-chiral limit cases. In the former case, the dominant excitations in the hadronic phase are the massless pions, while in the QGP phase are massless $u$, $d$ quarks and gluons. In the latter case, in the QGP phase the masses of $u$, $d$ and $s$ quarks are considered, and in the hadronic phase the hadrons and hadron resonances composed of $u$, $d$ and $s$ quarks with masses below $2.5$ GeV$/c^{2}$ are considered. Then we investigate the QCD phase transition with finite chemical potential in the non-chiral limit case. We will investigate the effect of the chemical potential and the non-extensive statistics on the temperature dependence of the dimensionless thermodynamic quantities such as the entropy density, the pressure, the energy density, the trace anomaly, the heat capacity at the constant volume, and the sound velocity squared. Moreover, we will compare our results with those from the HRG, the LQCD, and the fit of the experimental yields of hadrons at the LHC and RHIC energies using the Thermal-Fist package [58].
@@
This paper is organized as follows. In sect. II, we will describe the Tsallis statistics and construct a non-extensive version of EoS by applying the non-extensive correction to the EoS in the parton (hadron resonance) gas at high (low) temperature and making a smooth interpolation of these two equation of states. In sect. III, we will discuss the effect of the non-extensive statistics and the chemical potential on the thermodynamic quantities and compare our results with different models. Finally, the conclusion is given in sect. IV.
@@ II.1 Tsallis statistics
where $\alpha$, $\beta$ and $\gamma$ are the Lagrange multipliers associated with the total probability, the total energy and the total number of particles, respectively. Following the standard thermodynamic definitions, such as $\beta=1/T$ and $\gamma=-\beta\mu$ with $T$ and $\mu$, respectively, being the temperature and the chemical potential of the system, the probability distribution of the state $R$ in the grand-canonical ensemble is written as
@@
where $\varepsilon_{k}$ and $n_{k}$ are, respectively, the energy of a particle and the number of particles in the state $k$. Substituting Eqs. (9) and (10) into Eq. (8), one gets the partition function as
@@
where $\textrm{exp}_{q}(x)=[1+(1-q)x]^{(1/(1-q))}$, $\textrm{exp}_{q}(x)\otimes_{q}\textrm{exp}_{q}(y)=\textrm{exp}_{q}(x+y)$, $(\prod)_{q}$ means the successive $q$-multiplication. When taking into account the $d$ internal degrees of freedom for the non-interacting bosons or fermions, the grand-canonical partition function becomes
$Z_{q}\approx(\prod_{k=1}^{\infty})_{q}\left[1\pm{\rm exp}_{q}\left(-\frac{\varepsilon_{k}-\mu}{T}\right)\right]^{\pm d},$ (12)
where the plus and minus signs, respectively, refer to fermions and bosons. The derivation of Eq. (12) is shown in the appendix A.
@@ II.2 Equation of states in the hadronic and QGP phases
where $P_{q}^{i}$ is the pressure for the single-component system with the particle type $i$. One key factor in the calculation of the entropy density is the internal degrees of freedom $d$. In the chiral limit case, the dominant excitations in the hadronic (QGP) phase are the massless pions ($u$, $d$ quarks and gluons). The degeneracy factor for massless pions is $d_{\pi}=N_{f}^{2}-1$, where $N_{f}=$ 2 is the number of quark flavors considered. The degeneracy factors for the massless quarks and gluons are, respectively, $d_{g}=2_{\textrm{spin}}\times(N_{c}^{2}-1)=$ 16 and $d_{q}=2_{\textrm{spin}}\times 2_{q\bar{q}}\times N_{c}\times N_{f}=$ 24, where $N_{c}=3$ is the number of colors. For the non-chiral limit case, we consider the masses of $u$, $d$ and $s$ quarks (hadrons composed of these quarks) in the QGP (hadronic) phase. The degeneracy factors for quarks (antiquarks) with a given species and gluons are, respectively, $d_{q(\bar{q})}=6$ and $d_{g}=16$, while the degeneracy factor for hadrons with a given species is $d_{H}=2S+1$, where $S$ is the hadron spin. Another key factor in the calculation of the entropy density is the chemical potential. The chemical potential for $i^{th}$ particle species is $\mu_{i}=B_{i}\mu_{B}+S_{i}\mu_{S}+Q_{i}\mu_{Q}$, where $B_{i}$, $S_{i}$ and $Q_{i}$ are, respectively, the particle’s baryon, strangeness and electric charge numbers, $\mu_{B}$, $\mu_{S}$ and $\mu_{Q}$ are the respective chemical potentials of the system.
@@ II.3 Smooth crossover around the critical temperature
In order to match the equation of states in the hadronic and QGP phases around the crossover temperature $T_{c}$, we make a smooth interpolation of $s_{q}(T)$ between the hadronic gas at low $T$ and the QGP at high $T$. A simplest possible parametrization of the entropy density $s_{q}(T)$ satisfying the thermodynamic inequality $\partial s_{q}(T)/\partial T\geq 0$ and the third law of thermodynamics $s_{q}(T\rightarrow 0)\rightarrow\textrm{constant}$ is [12]
@@
The pressure $P_{q}(T)$, the energy density $\varepsilon_{q}(T)$, the specific heat at constant volume $(C_{V})_{q}(T)$, and the sound velocity squared $(c_{s}^{2})_{q}(T)$ then can be, respectively, obtained using the following expressions,
[…]
@@ III.1 Zero chemical potential
For the system with zero chemical potential, we will first consider the QCD phase transition in the chiral limit case, where the dominant excitations in the hadronic (QGP) phase are the massless pions (massless $u$, $d$ quarks and gluons). In order to see the mass effect, we also consider the QCD phase transition in the non-chiral limit case. In this case, the $u$, $d$ and $s$ quarks in the QGP phase are considered, and their masses are, respectively, $m_{u}=3.503$ MeV$/c^{2}$, $m_{d}=3.503$ MeV$/c^{2}$, and $m_{s}=96.4$ MeV$/c^{2}$ [62]. In the hadronic phase, hadrons and hadron resonances composed of $u$, $d$ and $s$ quarks with masses below $2.5$ GeV$/c^{2}$ [16] from the Particle Data Group [63] are considered.
Figure 1: The upper (lower) panels: $s_{q}/T^{3}$ $(P_{q}/T^{4})$ as a function of $T/T_{c}$. The left (right) panels correspond to the chiral (non-chiral) limit case. The solid and dashed lines correspond to $q=$ 1.05 and $q=$ 1, respectively.
Figure 2: The upper (lower) panels: $\varepsilon_{q}/T^{4}$ $((C_{V})_{q}/T^{3})$ as a function of $T/T_{c}$. The left (right) panels correspond to the chiral (non-chiral) limit case. The solid and dashed lines correspond to $q=$ 1.05 and $q=$1, respectively.
Figure 3: The upper (lower) panels: $\Delta_{q}$ ($(c_{s}^{2})_{q}$) as a function of $T/T_{c}$. The left (right) panels correspond to the chiral (non-chiral) limit case. The solid and dashed black lines correspond to $q=$1.05 and $q=$ 1, respectively. In the lower left panel, the solid and dashed blue lines correspond to $m_{\pi}$= 140 ${\rm MeV}/c^{2}$.
Figures 1, 2 and 3, respectively, present the dimensionless thermodynamic quantities such as the entropy density $s_{q}/T^{3}$, the pressure $P_{q}/T^{4}$, the energy density $\varepsilon_{q}/T^{4}$, the specific heat at constant volume $(C_{V})_{q}/T^{3}$, the trace anomaly $\Delta_{q}=(\varepsilon_{q}-3P_{q})/T^{4}$ and the sound velocity squared $(c_{s}^{2})_{q}$ as a function of the temperature for two non-extensive parameters $q=$ 1 (the dashed line) and $q=$ 1.05 (the solid line). The left (right) panels of these figures show the temperature dependence of these thermodynamic quantities in the chiral (non-chiral) limit case. In these figures, $T_{c}$ is taken to be 156.5 MeV [64], $\it\Gamma$ is set as 0.05 $T_{c}$, which is identical to the choice made in Ref. [12]. The qualitative behavior of these thermodynamic quantities in both cases can be summarized as follows.
(i) There is a rapid increase of the entropy density in a narrow region of $T$ around $T_{c}$ (with a width $\sim$ 8 MeV). In the high temperature region, $s_{q}/T^{3}$ approaches the Stefan-Boltzmann (SB) limit, and the SB value in the chiral limit case is smaller than that in non-chiral limit case. This is due to reason that in the latter case one additional quark flavor is considered. In the low temperature region, $s_{q}/T^{3}$ approaches a constant non-zero value for the former case, while it approaches 0 for the latter case.
@@
(v) $\Delta_{q}$ also has a prominent peak near $T_{c}$. Such a peak is a consequence of the rapid increase of $s_{q}(T)$. It vanishes when the temperature approaches 0 and $\infty$.
(vi) $s_{q}/T^{3}$, $P_{q}/T^{4}$, $\varepsilon_{q}/T^{4}$, $(C_{V})_{q}/T^{3}$ and $\Delta_{q}$ are sensitive to the deviation of $q$ from unity and they become large both in the hadronic and QGP phases with the increase of $q$. This can be interpreted as the effect that at a fixed temperature the particle yield in the non-extensive distribution with $q>$ 1 is always larger than that in the extensive one.
(vii) There is a sudden drop of $(c_{s}^{2})_{q}$ in the narrow region $|T-T_{c}|<\it\Gamma$. This could be explained as follows. $P_{q}(T)$ $(\varepsilon_{q}(T))$ varies slowly (rapidly) with temperature across the critical point, which leads to the softening of the equation of state in the critical region. In order to see the effect of the finite pion mass on $(c_{s}^{2})_{q}$ below $T_{c}$, the hadroinc sound velocity with $m_{\pi}=$ 140 ${\rm MeV}/c^{2}$ (blue solid and dashed lines) is also presented in the lower left panel of Fig. 3. In the chiral limit case, away from $T_{c}$, $(c_{s}^{2})_{q}$ approaches the relativistic limit, $1/3$. In the non-chiral limit case, $(c_{s}^{2})_{q}$ approaches $1/3$ in the high temperature region and 0 in the low temperature region. Moreover, in the former case, $(c_{s}^{2})_{q}$ is insensitive to the deviation of $q$ from unity. However, in the latter case, this deviation leads to nontrivial corrections of the sound velocity in the vicinity of the critical point and at lower temperatures.

### v2

@@ Abstract
The intrinsic fluctuations, memory effects and long-range color interactions in high energy nuclear collisions imply the presence of non-Markovian processes in the fireball evolution, which affects the thermalization process towards equilibrium and produces a non-extensive behavior. In order to investigate the non-equilibrium effect on the quantum chromodynamics (QCD) phase transition at finite temperature ($T$) and chemical potential ($\mu$), we apply a non-extensive correction to the equation of state in the parton (hadron resonance) gas at high (low) temperature and interpolate these two equation of states with a smooth crossover. The non-extensive statistics is characterized by a non-extensivity parameter $q$, which measures the degrees of deviation from the thermal equilibrium. It is found that the dimensionless thermodynamic quantities such as the entropy density, the pressure, the energy density, the specific heat at constant volume and the trace anomaly are sensitive to the deviation of $q$ from unity and they become large both in the hadronic and quark-gluon plasma phases with the increase of $q$. Moreover, this deviation leads to nontrivial corrections of the squared speed of sound ($(c_{s}^{2})_{q}$) in the vicinity of the critical point ($T_{c}$) and at lower temperatures. Additionally, these thermodynamic quantities are sensitive to the deviation of $\mu$ from zero. With increasing $\mu$, they become enhanced in both phases. Specifically, for $(c_{s}^{2})_{q}$, the value increases near $T_{c}$ but decreases at lower temperatures. Finally, we observe that our results with $q=1$ agree well with those from the Lattice QCD, the hadron resonance gas model, and the Thermal-Fist fit to the hadron yields in high energy nuclear collisions in the low temperature region up to $T\sim 150$ MeV.
@@ I Introduction
Several effective models, such as the bag model [1], the Nambu-Jona-Lasinio (NJL) model [2, 3], the linear sigma model [4], the quark-meson model [5, 6], the Dyson-Schwinger equations approach [7, 8], the functional renormalization group approach [9, 10, 11], the hadron resonance gas (HRG) and statistical models [12, 13, 14, 15], and the Lattice QCD (LQCD) [14, 15, 16, 17], have been utilized to investigate the properties of the QCD phase transition. In these models, a thermal equilibrium is assumed and the extensive statistical approach, such as the Boltzmann-Gibbs (BG) statistics, the Bose-Einstein statistics or the Fermi-Dirac statistics is applied.
@@
However, such an approach is only valid when the heat bath of the system is homogeneous. In reality, this condition cannot always be fulfilled. For example, in high-energy nucleus-nucleus (AA) collisions, extreme conditions of density and temperature give rise to strong intrinsic fluctuations, memory effects and long-range correlations [18, 19, 20, 21]. These conditions imply the presence of non-Markovian processes in the fireball evolution, which affects the thermalization process towards equilibrium and the standard equilibrium distribution [22]. Thus, instead of an equilibrium state, some kind of stationary state near the equilibrium ($q$-equilibrium) is expected to be formed [23, 24]. This phenomenon can be dealt with non-extensive statistics (Tsallis statistics) [25] where the degree of deviation from the equilibrium is described by a single parameter $q$. The non-extensivity in a QCD system can also be understood as a result of the fractal structure of the hot and dense system formed in high energy nuclear collisions [26, 27].
@@
In recent years, Tsallis statistics has been widely employed to describe the multi-particle production processes in high-energy proton-proton (pp) [28, 29, 30, 31, 32, 33, 34, 35, 36] and AA [37, 38, 39, 40, 41, 42, 43, 44, 45] collisions. In most of these applications, the Tsallis distribution or the Tsallis-extended blast-wave model is simply fitted to the particle transverse momentum ($p_{\rm T}$) spectra. The extracted values of $q$ from the fits are usually larger than 1. The deviation of $q$ from unity represents the intrinsic fluctuations of the temperature [18] or of the mean value of the charged-particle multiplicity [46] in the hadronizing system. This deviation can be also be related to the fractal dimension of the hadronizing system [26, 27]. In case of high-energy AA collisions, the fluid dynamics is not considered in these applications, which could affect the accuracy of the results. In light of such a situation, (0+1)-, (1+1)-, (2+1)-, and (3+1)-dimensional hydrodynamic models based on non-extensive statistics were developed as a description of relativistic heavy-ion collisions [47, 48, 49, 50]. This is a reasonable extension of the hydrodynamic model, as far-from-equilibrium evolution can still obey the hydrodynamic equations of motion, provided that the system is locally isotropized in the co-moving frame of the fluid [51, 52]. It was shown that the non-extensive hydrodynamic model can describe the charged hadron $p_{\rm T}$ spectra up to 6-8 GeV/c, which is larger than the typical validity range of the extensive hydrodynamic description, 2-3 GeV/c. Moreover, the elliptic flow of charged particles was reduced with the introduction of Tsallis statistics to the hydrodynamic model. The Tsallis statistics has also been applied on the nuclear equation of state (EoS) and the QGP EoS in the framework of a relativistic mean-field approach and the bag model, respectively. In the former model, the interactions of nucleons are mediated by the exchange of isoscalar and isovector mesons, $\sigma$, $\omega$, and $\rho$ [53, 54, 55, 56]. In the latter model, the quark matter is described as a gas of free quarks and the non-perturbative effects are simulated by the bag constant $B$ [53, 54, 56]. It was observed that both the nuclear EoS and the QGP EoS became stiffer when the non-extensive statistic effect was considered. Furthermore, the Tsallis statistics has been taken into account in the NJL model to describe the QCD phase transition [57, 58, 59, 60]. It was found that with the increase of $q$ the critical end point moved toward the direction of larger chemical potential and lower temperature.
[…]
@@
This paper is organized as follows. In sect. II, we will describe the Tsallis statistics and construct a non-extensive version of EoS by applying the non-extensive correction to the EoS in the parton (hadron resonance) gas at high (low) temperature and making a smooth interpolation of these two equation of states. In sect. III, we will discuss the effects of the non-extensive statistics and the chemical potential on the thermodynamic quantities and compare our results with different models. Finally, the conclusion is given in sect. IV.
@@ II.1 Tsallis statistics
where $\alpha$, $\beta$ and $\gamma$ are the Lagrange multipliers associated with the total probability, the total energy, and the total number of particles, respectively. Following the standard thermodynamic definitions, such as $\beta=1/T$ and $\gamma=-\beta\mu$ with $T$ and $\mu$, respectively, being the temperature and the chemical potential of the system, the probability distribution of the state $R$ in the grand-canonical ensemble is written as
@@
where $\varepsilon_{k}$ and $n_{k}$ are, respectively, the energy of a particle and the number of particles in the state $k$. Substituting Eqs. (9) and (10) into Eq. (8), one gets the partition function as [65]
@@
where $\textrm{exp}_{q}(x)=[1+(1-q)x]^{(1/(1-q))}$, $\textrm{exp}_{q}(x)\otimes_{q}\textrm{exp}_{q}(y)=\textrm{exp}_{q}(x+y)$, $(\prod)_{q}$ means the successive $q$-multiplication. When taking into account the $d$ internal degrees of freedom for the non-interacting fermions and bosons, the grand-canonical partition function, respectively, becomes
$Z_{q}=(\prod_{k=1}^{\infty})_{q}\left[1+{\rm exp}_{q}\left(-\frac{\varepsilon_{k}-\mu}{T}\right)\right]^{+d}$ (12)
and
$Z_{q}\approx(\prod_{k=1}^{\infty})_{q}\left[1-{\rm exp}_{q}\left(-\frac{\varepsilon_{k}-\mu}{T}\right)\right]^{-d}.$ (13)
The derivations of Eqs. (12) and (13) are shown in the appendix A. In the derivation of Eq. (13), we have made an approximation that the higher orders $\mathcal{O}(q-1)^{2}$ have been ignored.
@@ II.2 Equation of states in the hadronic and QGP phases
where $P_{q}^{i}$ is the pressure for the single-component system with the particle type $i$. One key factor in the calculation of the entropy density is the internal degrees of freedom $d$. In the QGP (hadronic) phase, the masses of $u$, $d$, and $s$ quarks (hadrons and resonances composed of these quarks) are considered. The degeneracy factor for quarks (antiquarks) of a given flavor is $d_{q(\bar{q})}=2_{\text{spin}}\times N_{c}=6$, while for gluons the degeneracy is $d_{g}=2_{\text{spin}}\times(N_{c}^{2}-1)=16$, where $N_{c}=3$ represents the number of colors. For hadrons of a specific species, the degeneracy factor is given by $d_{H}=2S+1$, with $S$ denoting the spin of the hadron. Another key factor in the calculation of the entropy density is the chemical potential. The chemical potential for the $i^{th}$ particle species is $\mu_{i}=B_{i}\mu_{B}+S_{i}\mu_{S}+Q_{i}\mu_{Q}$, where $B_{i}$, $S_{i}$, and $Q_{i}$ are, respectively, the particle’s baryon, strangeness, and electric charge numbers, $\mu_{B}$, $\mu_{S}$, and $\mu_{Q}$ are the respective chemical potentials of the system.
@@ II.3 Smooth crossover around the critical temperature
In section II.2, the EoS in the hadronic (QGP) phase has been constructed by applying a non-extensive correction to the hadron resonance (parton) gas model. In order to match the equation of states in the hadronic and QGP phases around the crossover temperature $T_{c}$, we make a smooth interpolation of $s_{q}(T)$ between the hadronic gas at low $T$ and the QGP at high $T$. A simplest possible parametrization of the entropy density $s_{q}(T)$ satisfying the thermodynamic inequality $\partial s_{q}(T)/\partial T\geq 0$ and the third law of thermodynamics $s_{q}(T\rightarrow 0)\rightarrow\textrm{constant}$ is [12]
@@
The pressure $P_{q}(T)$, the energy density $\varepsilon_{q}(T)$, the specific heat at constant volume $(C_{V})_{q}(T)$, and the squared speed of sound $(c_{s}^{2})_{q}(T)$ then can be, respectively, obtained using the following expressions,
$P_{q}(T)=\int_{0}^{T}s_{q}(t)dt,$ (23)
$\varepsilon_{q}(T)=Ts_{q}(T)-P_{q}(T)+\sum_{i}\mu_{i}\int_{0}^{T}\frac{\partial s_{q}(t)}{\partial\mu_{i}}dt,$ (24)
@@ III.1 Zero chemical potential
[…]
In each subplot of Fig. 1, the dashed and solid back curves, respectively, present the dimensionless thermodynamic quantities such as the entropy density $s_{q}/T^{3}$, the pressure $P_{q}/T^{4}$, the energy density $\varepsilon_{q}/T^{4}$, the specific heat at constant volume $(C_{V})_{q}/T^{3}$, the trace anomaly $\Delta_{q}=(\varepsilon_{q}-3P_{q})/T^{4}$, and the squared speed of sound $(c_{s}^{2})_{q}$ as a function of the temperature for two non-extensive parameters $q=$ 1 and $q=$ 1.05. In the figure, $T_{c}$ is taken to be 156.5 MeV [70], $\it\Gamma$ is set as 0.05 $T_{c}$, which is identical to the choice made in Ref. [12]. The qualitative behavior of these thermodynamic quantities in both cases can be summarized as follows.
Figure 1: (a) ((b), (c), (d), (e), and (f)): $s_{q}/T^{3}$ ($P_{q}/T^{4}$, $\varepsilon_{q}/T^{4}$, $\Delta_{q}$, $(C_{V})_{q}/T^{3}$, and $(c_{s}^{2})_{q}$) as a function of $T/T_{c}$. The solid and dashed black lines correspond to $q=$ 1.05 and $q=$ 1 with zero chemical potential, respectively. The solid and dashed blue lines correspond to $q=$ 1.05 and $q=$ 1 with $\mu_{B}=$ 398.2 MeV, $\mu_{S}=$ 89.5 MeV, and $\mu_{Q}=$ 0 MeV, respectively.
(i) There is a rapid increase of the entropy density in a narrow region of $T$ around $T_{c}$ (with a width $\sim$ 8 MeV). In the high temperature region, $s_{q}/T^{3}$ approaches the Stefan-Boltzmann (SB) limit. In the low temperature region, $s_{q}/T^{3}$ approaches 0. This behavior could be explained as follows. The excitation of a massive hadron of mass $m$ is exponentially suppressed by the Boltzmann factor $\exp(-m/T)$. The entropy density $s_{q}$ is dominated by this exponential factor, which vanishes much more rapidly than the power-law term $T^{3}$ when $T\rightarrow 0$.
@@
(v) $\Delta_{q}$ also has a prominent peak near $T_{c}$. Such a peak is a consequence of the rapid increase of $s_{q}(T)$. It vanishes when the temperature approaches either0 or $\infty$.
(vi) $s_{q}/T^{3}$, $P_{q}/T^{4}$, $\varepsilon_{q}/T^{4}$, $(C_{V})_{q}/T^{3}$, and $\Delta_{q}$ are sensitive to the deviation of $q$ from unity and they become large both in the hadronic and QGP phases with the increase of $q$. This can be interpreted as the effect that at a fixed temperature the particle yield in the non-extensive distribution with $q>$ 1 is always larger than that in the extensive one.
Figure 2: (a) ((b), (c), (d), (e), and (f)): $s_{q}/T^{3}$ ($P_{q}/T^{4}$, $\varepsilon_{q}/T^{4}$, $\Delta_{q}$, $(C_{V})_{q}/T^{3}$, and $(c_{s}^{2})_{q}$) as a function of $T/T_{c}$ at zero chemical potential. The dashed (solid) black and blue curves, respectively, represents the temperature dependence of these quantities with $\it\Gamma=\textrm{0.05 }T_{c}$ and $\it\Gamma=\textrm{0.25 }T_{c}$ for the extensive (non-extensive) case.
(vii) $(c_{s}^{2})_{q}$ exhibits a sharp drop in the critical region $|T-T_{c}|<\it\Gamma$ and stays below 1/3 in the hadronic phase. In both cases, the behavior results from the energy density $\varepsilon_{q}(T)$ rising more rapidly than the pressure $P_{q}(T)$, leading to a soft equation of state. Across $T_{c}$, this is due to deconfinement. In the hadronic phase, it is a consequence of hadron masses. Moreover, $(c_{s}^{2})_{q}$ approaches $1/3$ in the high temperature region and 0 in the low temperature region. This could be understood as follows. At low temperatures, the exponential suppression $\exp(-m/T)$ causes both $P_{q}$ and $\varepsilon_{q}$ to vanish, but their derivatives vanish at different rates. The ratio of these derivatives, $(c_{s}^{2})_{q}=\partial P_{q}/\partial\varepsilon_{q}$, scales as $T/m$, guaranteeing it approaches zero with $T\rightarrow 0$. At extremely high temperatures, the masses of the $u$, $d$, and $s$ are negligible. Thus, these quarks behave dynamically as massless non-interacting particles. Additionally, the deviation of $q$ from unity leads to nontrivial corrections of the squared speed of sound in the vicinity of the critical point and at lower temperatures.
We have also investigated the $\it\Gamma$ dependence of the thermodynamic quantities at the vanishing chemical potential for both extensive and non-extensive cases in Fig. 2. The dashed (solid) black and blue curves, respectively, represents the temperature dependence of these quantities with $\it\Gamma=\textrm{0.05 }T_{c}$ and $\it\Gamma=\textrm{0.25 }T_{c}$ for the extensive (non-extensive) case. It is observed that an increase in the value of $\it\Gamma$ leads to a more gradual temperature dependence of the entropy density and pressure. Concurrently, the characteristic peaks in the trace anomaly and the specific heat at constant volume undergo a broadening and a reduction in height. For the squared speed of sound, an increase in $\it\Gamma$ produces a dip that is both shallower and significantly wider. This set of phenomena is attributable to the broadening of the QCD phase transition window. Specifically, a larger $\it\Gamma$ implies that the conversion from the hadronic phase to the QGP phase takes place across an extended temperature interval, which smears out the thermodynamic profiles, resulting in smoother, broader, and flatter curves.